OpenStudy (anonymous):

a football player kicks a ball at an angle of 36 degrees above the ground with an initial velocity of 60 feet per second. write the parametric equations for the position of the football at time t seconds. use g=32ft/sec^2.

OpenStudy (anonymous):

I got x(t) = (60*cos(36)t ; y(t)= (60*sin(36))t-16t^2 but part (b) is asking me for the path that the football follows idk how to do this part

OpenStudy (mos1635):

by eleminate t from your equations go to find the eq y=f(x)

OpenStudy (ujjwal):

@Ldaniel You've mentioned two equations.. from the first one get the value of t in terms of x, and then replace it in 2nd eqn.. the newly formed equation is what you need..

OpenStudy (anonymous):

 X=vcos36*t.....(i) Y=vsin36*t-1/2gt^2...(ii) these two are parametric equation and eliminate t in both equation you will get equation of parabola.....